OpenStudy (anonymous):

Solve using the elimination method. If the system has no solution or an infinite number of solutions, state this. -2x + 6y = -12, x + 3y = -3

OpenStudy (anonymous):

-2x+6y=-12 - 2x+6y=-6 becomes -4x = -6 x =6/4 y = -9/6 Do you see how I did this? Wait that's what you meant by elimination right?

OpenStudy (anonymous):

Thanks! I'll have to work on this some more to get a better understanding.